Subfig Package Test Program One

Steven Douglas Cochran Digital Mapping Laboratory School of Computer Science Carnegie-Mellon University 5000 Forbes Avenue Pittsburgh, PA 15213-3890 USA

21 December 2003

Abstract

This test checks five things:

1. if the subfig and the fixltx2e packages work together;

2. that the subcaption label format for the List-of-Float pages is working;

3. that the subcaption numbering in the List-of-Float pages is working;

4. it tests the figure* and table* environemnts in two column mode; and,

5. that various position combinations of top and bottom captions and subcaptions work together.

List of Figures

1 First figure — but this one has a very very long caption. So long that it continues over into a second line so that we can test the margin setting and centering of the caption command in the full page mode. . . . . . . . . . . . . . 5 (a) One subone. . . . . . . . . . . . 5 (b) One subtwo. — but this one has

a very very long caption. So long that it continues over into other lines so that we can test the listof line settings. . . . . . . . . . . 5

2 Second figure. . . . . . . . . . . . . . . 5 (a) Two subone. . . . . . . . . . . . 5 (b) Two subtwo. . . . . . . . . . . . 5

3 Third figure. . . . . . . . . . . . . . . 5 (a) Three subone. . . . . . . . . . . 5 (b) Three subtwo. . . . . . . . . . . 5

4 Fourth figure. . . . . . . . . . . . . . . 6 (a) Four subone. . . . . . . . . . . . 6 (b) Four subtwo. . . . . . . . . . . . 6

5 Fifth figure. . . . . . . . . . . . . . . . 6 (a) Five subone. . . . . . . . . . . . 6 (b) Five subtwo. . . . . . . . . . . . 6

6 Sixth figure. . . . . . . . . . . . . . . . 6 (a) Six subone. . . . . . . . . . . . . 6 (b) Six subtwo. . . . . . . . . . . . . 6

7 First figure. . . . . . . . . . . . . . . . 9 (a) One subone. . . . . . . . . . . . 9 (b) One subtwo. . . . . . . . . . . . 9

8 Second figure. . . . . . . . . . . . . . . 9 (a) Two subone. . . . . . . . . . . . 9 (b) Two subtwo. . . . . . . . . . . . 9

9 Third figure. . . . . . . . . . . . . . . 9 (a) Three subone. . . . . . . . . . . 9 (b) Three subtwo. . . . . . . . . . . 9

10 Fourth figure. . . . . . . . . . . . . . . 10 (a) Four subone. . . . . . . . . . . . 10 (b) Four subtwo. . . . . . . . . . . . 10

11 Fifth figure. . . . . . . . . . . . . . . . 10 (a) Five subone. . . . . . . . . . . . 10 (b) Five subtwo. . . . . . . . . . . . 10

12 Sixth figure. . . . . . . . . . . . . . . . 10 (a) Six subone. . . . . . . . . . . . . 10

(b) Six subtwo. . . . . . . . . . . . . 10 13 First figure. . . . . . . . . . . . . . . . 13 (a) One subone. . . . . . . . . . . . 13 (b) One subtwo. . . . . . . . . . . . 13 14 Second figure. . . . . . . . . . . . . . . 13 (a) Two subone. . . . . . . . . . . . 13 (b) Two subtwo. . . . . . . . . . . . 13 15 Third figure. . . . . . . . . . . . . . . 13 (a) Three subone. . . . . . . . . . . 13 (b) Three subtwo. . . . . . . . . . . 13 16 Fourth figure. . . . . . . . . . . . . . . 14 (a) Four subone. . . . . . . . . . . . 14 (b) Four subtwo. . . . . . . . . . . . 14 17 Fifth figure. . . . . . . . . . . . . . . . 14 (a) Five subone. . . . . . . . . . . . 14 (b) Five subtwo. . . . . . . . . . . . 14 18 Sixth figure. . . . . . . . . . . . . . . . 14 (a) Six subone. . . . . . . . . . . . . 14 (b) Six subtwo. . . . . . . . . . . . . 14 19 First figure. . . . . . . . . . . . . . . . 17 (a) One subone. . . . . . . . . . . . 17 (b) One subtwo. . . . . . . . . . . . 17 20 Second figure. . . . . . . . . . . . . . . 17 (a) Two subone. . . . . . . . . . . . 17 (b) Two subtwo. . . . . . . . . . . . 17 21 Third figure. . . . . . . . . . . . . . . 18 (a) Three subone. . . . . . . . . . . 18 (b) Three subtwo. . . . . . . . . . . 18 22 First figure. . . . . . . . . . . . . . . . 19 (a) One subone. . . . . . . . . . . . 19 (b) One subtwo. . . . . . . . . . . . 19 23 Second figure. . . . . . . . . . . . . . . 19 (a) Two subone. . . . . . . . . . . . 19 (b) Two subtwo. . . . . . . . . . . . 19 24 Third figure. . . . . . . . . . . . . . . 20 (a) Three subone. . . . . . . . . . . 20 (b) Three subtwo. . . . . . . . . . . 20

List of Tables

1 First table. . . . . . . . . . . . . . . . 7 a: One subone. . . . . . . . . . . . . . 7 b: One subtwo. . . . . . . . . . . . . . 7

2 Second table. . . . . . . . . . . . . . . 7 a: Two subone. . . . . . . . . . . . . . 7 b: Two subtwo. . . . . . . . . . . . . 7

3 Third table. . . . . . . . . . . . . . . . 7 a: Three subone. . . . . . . . . . . . . 7 b: Three subtwo. . . . . . . . . . . . . 7

4 Fourth table. . . . . . . . . . . . . . . 8 a: Four subone. . . . . . . . . . . . . 8 b: Four subtwo. . . . . . . . . . . . . 8

5 Fifth table. . . . . . . . . . . . . . . . 8 a: Five subone. . . . . . . . . . . . . . 8 b: Five subtwo. . . . . . . . . . . . . . 8

6 Sixth table. . . . . . . . . . . . . . . . 8 a: Six subone. . . . . . . . . . . . . . 8 b: Six subtwo. . . . . . . . . . . . . . 8

7 First table. . . . . . . . . . . . . . . . 11 a: One subone. . . . . . . . . . . . . . 11 b: One subtwo. . . . . . . . . . . . . . 11

8 Second table. . . . . . . . . . . . . . . 11 a: Two subone. . . . . . . . . . . . . . 11 b: Two subtwo. . . . . . . . . . . . . 11

9 Third table. . . . . . . . . . . . . . . . 11 a: Three subone. . . . . . . . . . . . . 11 b: Three subtwo. . . . . . . . . . . . . 11

10 Fourth table. . . . . . . . . . . . . . . 12 a: Four subone. . . . . . . . . . . . . 12 b: Four subtwo. . . . . . . . . . . . . 12

11 Fifth table. . . . . . . . . . . . . . . . 12 a: Five subone. . . . . . . . . . . . . . 12 b: Five subtwo. . . . . . . . . . . . . . 12

12 Sixth table. . . . . . . . . . . . . . . . 12 a: Six subone. . . . . . . . . . . . . . 12 b: Six subtwo. . . . . . . . . . . . . . 12

13 First table. . . . . . . . . . . . . . . . 15 a: One subone. . . . . . . . . . . . . . 15 b: One subtwo. . . . . . . . . . . . . . 15

14 Second table. . . . . . . . . . . . . . . 15 a: Two subone. . . . . . . . . . . . . . 15 b: Two subtwo. . . . . . . . . . . . . 15

15 Third table. . . . . . . . . . . . . . . . 15 a: Three subone. . . . . . . . . . . . . 15

b: Three subtwo. . . . . . . . . . . . . 15 16 Fourth table. . . . . . . . . . . . . . . 16 a: Four subone. . . . . . . . . . . . . 16 b: Four subtwo. . . . . . . . . . . . . 16 17 Fifth table. . . . . . . . . . . . . . . . 16 a: Five subone. . . . . . . . . . . . . . 16 b: Five subtwo. . . . . . . . . . . . . . 16 18 Sixth table. . . . . . . . . . . . . . . . 16 a: Six subone. . . . . . . . . . . . . . 16 b: Six subtwo. . . . . . . . . . . . . . 16 19 First table. . . . . . . . . . . . . . . . 17 a: One subone. . . . . . . . . . . . . . 17 b: One subtwo. . . . . . . . . . . . . . 17 20 Second table. . . . . . . . . . . . . . . 18 a: Two subone. . . . . . . . . . . . . . 18 b: Two subtwo. . . . . . . . . . . . . 18 21 Third table. . . . . . . . . . . . . . . . 18 a: Three subone. . . . . . . . . . . . . 18 b: Three subtwo. . . . . . . . . . . . . 18 22 First table. . . . . . . . . . . . . . . . 19 a: One subone. . . . . . . . . . . . . . 19 b: One subtwo. . . . . . . . . . . . . . 19 23 Second table. . . . . . . . . . . . . . . 20 a: Two subone. . . . . . . . . . . . . . 20 b: Two subtwo. . . . . . . . . . . . . 20 24 Third table. . . . . . . . . . . . . . . . 20 a: Three subone. . . . . . . . . . . . . 20 b: Three subtwo. . . . . . . . . . . . . 20

1 Overview

This test runs through the various layout options for both figures and subfigures; and tables and subtables. It also shows a mixture of figures and tables within a single float environment.

1.1 Multiple figures

Figure 1[5] has two subfigures 1a[5] and 1b[5]. Figure 2[5] has two subfigures 2a[5] and 2b[5]. Figure 3[5] has two subfigures 3a[5] and 3b[5]. Figure 4[6] has two subfigures 4a[6] and 4b[6]. Figure 5[6] has two subfigures 5a[6] and 5b[6]. Figure 6[6] has two subfigures 6a[6] and 6b[6].

1.2 Multiple tables

Table 1[7] has two subtables 1a[7] and 1b[7]. Table 2[7] has two subtables 2a[7] and 2b[7]. Table 3[7] has two subtables 3a[7] and 3b[7]. Table 4[8] has two subtables 4a[8] and 4b[8]. Table 5[8] has two subtables 5a[8] and 5b[8]. Table 6[8] has two subtables 6a[8] and 6b[8].

1.3 Multiple figures with the figure caption at top

Figure 7[9] has two subfigures 7a[9] and 7b[9]. Figure 8[9] has two subfigures 8a[9] and 8b[9]. Figure 9[9] has two subfigures 9a[9] and 9b[9]. Figure 10[10] has two subfigures 10a[10] and 10b[10]. Figure 11[10] has two subfigures 11a[10] and 11b[10]. Figure 12[10] has two subfigures 12a[10] and 12b[10].

1.4 Multiple tables with the table caption at top

Table 7[11] has two subtables 7a[11] and 7b[11]. Table 8[11] has two subtables 8a[11] and 8b[11]. Table 9[11] has two subtables 9a[11] and 9b[11]. Table 10[12] has two subtables 10a[12] and 10b[12]. Table 11[12] has two subtables 11a[12] and 11b[12]. Table 12[12] has two subtables 12a[12] and 12b[12].

1.5 Multiple figures with the figure and subfigure captions at top

Figure 13[13] has two subfigures 13a[13] and 13b[13]. Figure 14[13] has two subfigures 14a[13] and 14b[13]. Figure 15[13] has two subfigures 15a[13] and 15b[13]. Figure 16[14] has two subfigures 16a[14] and 16b[14]. Figure 17[14] has two subfigures 17a[14] and 17b[14]. Figure 18[14] has two subfigures 18a[14] and 18b[14].

1.6 Multiple tables with the table and subtable captions at top

Table 13[15] has two subtables 13a[15] and 13b[15]. Table 14[15] has two subtables 14a[15] and 14b[15]. Table 15[15] has two subtables 15a[15] and 15b[15]. Table 16[16] has two subtables 16a[16] and 16b[16]. Table 17[16] has two subtables 17a[16] and 17b[16]. Table 18[16] has two subtables 18a[16] and 18b[16].

1.7 Mixed figures and tables with only the table caption at top

Figure 19[17] has two subfigures 19a[17] and 19b[17]. Table 19[17] has two subtables 19a[17] and 19b[17]. Figure 20[17] has two subfigures 20a[17] and 20b[17]. Figure 21[18] has two subfigures 21a[18] and 21b[18]. Table 20[18] has two subtables 20a[18] and 20b[18]. Table 21[18] has two subtables 21a[18] and 21b[18].

1.8 Mixed figures and tables with table and subtable captions at top

Figure 22[19] has two subfigures 22a[19] and 22b[19]. Table 22[19] has two subtables 22a[19] and 22b[19]. Figure 23[19] has two subfigures 23a[19] and 23b[19]. Figure 24[20] has two subfigures 24a[20] and 24b[20]. Table 23[20] has two subtables 23a[20] and 23b[20]. Table 24[20] has two subtables 24a[20] and 24b[20].

One subone

(a) One subone.

One subtwo

(b) One subtwo. — but this one has a very very long caption. So long that it continues over into other lines so that we can test the list-of line settings.

Figure 1: First figure — but this one has a very very long caption. So long that it continues over into a second line so that we can test the margin setting and centering of the caption command in the full page mode.

Two subone

Two subtwo

Figure 2: Second figure.

Three subone

Three subtwo

Figure 3: Third figure.

(a) Two subone.

(b) Two subtwo.

(a) Three subone.

(b) Three subtwo.

Four subone

(a) Four subone.

Four subtwo

(b) Four subtwo.

Figure 4: Fourth figure.

Five subone

(a) Five subone.

Five subtwo

(b) Five subtwo.

Figure 5: Fifth figure.

Six subone

(a) Six subone.

Six subtwo

(b) Six subtwo.

Figure 6: Sixth figure.

One subone

(a) One subone.

One subtwo

(b) One subtwo.

Table 1: First table.

Two subone

(a) Two subone.

Two subtwo

(b) Two subtwo.

Table 2: Second table.

Three subone

(a) Three subone.

Three subtwo

(b) Three subtwo.

Table 3: Third table.

Four subone

(a) Four subone.

Four subtwo

(b) Four subtwo.

Table 4: Fourth table.

Five subone

(a) Five subone.

Five subtwo

(b) Five subtwo.

Table 5: Fifth table.

Six subone

(a) Six subone.

Six subtwo

(b) Six subtwo.

Table 6: Sixth table.

Figure 7: First figure.

(a) One subone.

One subone

(b) One subtwo.

One subtwo

Figure 8: Second figure.

(a) Two subone.

Two subone

(b) Two subtwo.

Two subtwo

Figure 9: Third figure.

(a) Three subone.

Three subone

(b) Three subtwo.

Three subtwo

Figure 10: Fourth figure.

(a) Four subone.

Four subone

(b) Four subtwo.

Four subtwo

Figure 11: Fifth figure.

(a) Five subone.

Five subone

(b) Five subtwo.

Five subtwo

Figure 12: Sixth figure.

(a) Six subone.

Six subone

(b) Six subtwo.

Six subtwo

Table 7: First table.

(a) One subone.

One subone

(b) One subtwo.

One subtwo

Table 8: Second table.

(a) Two subone.

Two subone

(b) Two subtwo.

Two subtwo

Table 9: Third table.

(a) Three subone.

Three subone

(b) Three subtwo.

Three subtwo

Table 10: Fourth table.

(a) Four subone.

Four subone

(b) Four subtwo.

Four subtwo

Table 11: Fifth table.

(a) Five subone.

Five subone

(b) Five subtwo.

Five subtwo

Table 12: Sixth table.

(a) Six subone.

Six subone

(b) Six subtwo.

Six subtwo

Figure 13: First figure.

(a) One subone.

One subone

(b) One subtwo.

One subtwo

Figure 14: Second figure.

(a) Two subone.

Two subone

(b) Two subtwo.

Two subtwo

Figure 15: Third figure.

(a) Three subone.

Three subone

(b) Three subtwo.

Three subtwo

Figure 16: Fourth figure.

(a) Four subone.

Four subone

(b) Four subtwo.

Four subtwo

Figure 17: Fifth figure.

(a) Five subone.

Five subone

(b) Five subtwo.

Five subtwo

Figure 18: Sixth figure.

(a) Six subone.

Six subone

(b) Six subtwo.

Six subtwo

Table 13: First table.

(a) One subone.

One subone

(b) One subtwo.

One subtwo

Table 14: Second table.

(a) Two subone.

Two subone

(b) Two subtwo.

Two subtwo

Table 15: Third table.

(a) Three subone.

Three subone

(b) Three subtwo.

Three subtwo

Table 16: Fourth table.

(a) Four subone.

Four subone

(b) Four subtwo.

Four subtwo

Table 17: Fifth table.

(a) Five subone.

Five subone

(b) Five subtwo.

Five subtwo

Table 18: Sixth table.

(a) Six subone.

Six subone

(b) Six subtwo.

Six subtwo One subone

(a) One subone.

One subtwo

(b) One subtwo.

Figure 19: First figure. Table 19: First table.

One subone

(a) One subone.

One subtwo

(b) One subtwo.

Two subone

(a) Two subone.

Two subtwo

(b) Two subtwo.

Figure 20: Second figure.

Table 20: Second table.

Two subone

(a) Two subone.

Two subtwo

(b) Two subtwo.

Three subone

(a) Three subone.

Three subtwo

(b) Three subtwo.

Figure 21: Third figure. Table 21: Third table.

Three subone Three subtwo One subone

(a) Three subone.

(b) Three subtwo.

(a) One subone.

One subtwo

(b) One subtwo.

Figure 22: First figure. Table 22: First table.

(a) One subone.

One subone

(b) One subtwo.

One subtwo

Two subone

(a) Two subone.

Two subtwo

(b) Two subtwo.

Figure 23: Second figure.

Table 23: Second table.

(a) Two subone.

Two subone

(b) Two subtwo.

Two subtwo

Three subone

(a) Three subone.

Three subtwo

(b) Three subtwo.

Figure 24: Third figure. Table 24: Third table.

(a) Three subone.

Three subone

(b) Three subtwo.

Three subtwo